OpenStudy (anonymous):

solve for θ: csc^2θ=cotθ+1

OpenStudy (anonymous):

\[ \frac{1}{sina} = \frac{cosa + sina}{sina}\]

OpenStudy (anonymous):

No pi. pi in the cosine gives -1.

OpenStudy (anonymous):

ah i didn't see that it's cos^2

OpenStudy (anonymous):

\[\frac{1}{sin^2a} = \frac{cosa+sina}{sina}\] \[1 = sinacosa + sin^2a\] a = pi/2